# Loss Functions and Metrics in Deep Learning. A Review

### v1

@@ 4 Classification
Classification is a supervised machine learning task in which a model is trained to predict the class or category of a given input data point. Classification aims to learn a mapping from input features to a specific class or category.
@@
There are different classification tasks, such as binary classification, multi-class classification, and multi-label classification. Binary classification is a task where the model is trained to predict one of two classes, such as "spam" or "not spam," for an email. Multi-class classification is a task where the model is trained to predict one of the multiple classes, such as "dog," "cat," and "bird," for an image. Multi-label classification is a task where the model is trained to predict multiple labels for a single data point, such as "dog" and "outdoor," for an image of a dog in the park.
@@
Classification algorithms can be based on techniques such as decision trees, Naive Bayes, k-nearest neighbors, Support Vector Machines, Random Forest, Gradient Boosting, Neural Networks, and others.
4.1 Classification Loss Functions
@@
Several loss functions can be used for classification tasks, depending on the specific problem and algorithm. In the following sections, we describe the most common loss functions used for classification:
4.1.1 Binary Cross-Entropy Loss (BCE)
@@
The Binary Cross Entropy (BCE), also known as log loss, is a commonly used loss function for binary classification problems. It measures the dissimilarity between the predicted probability of a class and the true class label [57]. Cross-entropy is a well-known concept in information theory commonly used to measure the dissimilarity between two probability distributions. In binary classification, the true class is usually represented by a one-hot encoded vector, where the true class has a value of 1, and the other class has a value of 0. The predicted probability is represented by a vector of predicted probabilities for each class, where the predicted probability of the true class is denoted by $p(y=1|x)$ and the predicted probability of the other class is denoted by $p(y=0|x)$.
The loss function is defined as
@@
Which intuitively can be split into two parts:
@@
where $y$ is the true class label (0 or 1) and $p$ is the predicted probability of the positive class. The loss function is minimized when the predicted probability $p$ equals the true class label $y$.
The binary cross-entropy loss has several desirable properties, such as being easy to compute, differentiable, and providing a probabilistic interpretation of the model’s output. It also provides a smooth optimization surface and is less sensitive to outliers than other loss functions. However, it is sensitive to the class imbalance problem, which occurs when the number of samples of one class is significantly greater than the other. We can use the Weighted Binary Cross Entropy for these cases.
4.1.2 Weighted Binary Cross Entropy (WBCE)
Variation of the standard binary cross-entropy loss function, where the weight of each sample is considered during the loss calculation. This is useful in situations where the distribution of the samples is imbalanced [58].
In the standard binary cross-entropy loss, the loss is calculated as the negative log-likelihood of the true labels given the predicted probabilities. In the Weighted Binary Cross Entropy (WBCE), a weight is assigned to each sample, and the loss for each sample is calculated as
$L=-(w_{i}\cdot ylog(p)+w_{i}(1-y)log(1-p)),$ (18)
where $w_{i}$ is the weight assigned to the $i^{th}$ sample, $y$ is the true label, and $p$ is the predicted probability of the positive class.
By assigning a higher weight to samples from under-represented classes, the model is encouraged to pay more attention to these samples, and the model’s overall performance can be improved.
4.1.3 Categorical Cross-entropy Loss (CCE)
@@
The Categorical Cross Entropy (CCE), also known as the negative log-likelihood loss or Multi-class log loss, is a function used for multi-class classification tasks. It measures the dissimilarity between the predicted probability distribution and the true distribution [59].
@@
Given the predicted probability distribution, it is defined as the average negative log-likelihood of the true class. The formula for categorical cross-entropy loss is expressed as
$L=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{C}y_{i,j}log(p_{i,j}),$ (19)
where $N$ is the number of samples, $C$ is the number of classes, $y$ is the true label, and $p$ is the predicted probability of the true class. The loss is calculated for each sample and averaged over the entire dataset.
The true label is a one-hot encoded vector in traditional categorical cross-entropy loss, where the element corresponding to the true class is one, and all other elements are 0. However, in some cases, it is more convenient to represent the true class as an integer, where the integer value corresponds to the index of the true class leading to the sparse categorical cross-entropy loss discussed next.
4.1.4 Sparse Categorical Cross-entropy Loss
Variation of the categorical cross-entropy loss used for multi-class classification tasks where the classes are encoded as integers rather than one-hot encoded vectors [59]. Given that the true labels are provided as integers, we directly select the correct class using the provided label index instead of summing over all possible classes. Thus the loss for each example is calculated as
@@ 4.1.4 Sparse Categorical Cross-entropy Loss
And the final sparse categorical cross-entropy loss is the average over all the samples:
@@
where $y_{i}$ is the true class of the $i$-th sample and $\hat{y}_{i,y_{i}}$ is the predicted probability of the $i$-th sample for the correct class $y_{i}$.
@@
4.1.5 Cross-Entropy loss with label smoothing
In the Cross-Entropy loss with label smoothing, the labels are smoothed by adding a small value to the true label and subtracting the same value from all other labels. This helps reduce the model’s overconfidence by encouraging it to produce more uncertain predictions [60, 61].
The motivation behind this is that when training a model, it is common to become over-confident in its predictions, particularly when trained on a large amount of data. This overconfidence can lead to poor performance on unseen data. Label smoothing helps to mitigate this problem by encouraging the model to make less confident predictions.
The formula for the Cross-Entropy loss with label smoothing is similar to the standard categorical cross-entropy loss but with a small epsilon added to the true label and subtracted from all other labels. The formula is given by
$L(y,\hat{y})=-\sum_{c=1}^{C}\left[(1-\epsilon)y_{c}\log\hat{y}_{c}+\frac{\epsilon}{C}\log\hat{y}_{c}\right],$ (22)
where $y$ is the true label, $\hat{y}$ is the predicted label, $C$ is the number of classes, and $\epsilon$ is the smoothing value. Typically, $\epsilon$ is set to a small value, such as 0.1 or 0.2.
Label smoothing does not always improve performance, and it is common to experiment with different epsilon values to find the best value for a specific task and dataset.
4.1.6 Focal loss
The focal loss introduced by Tsung-Yi Lin et al. [62] is a variation of the standard cross-entropy loss that addresses the issue of class imbalance, which occurs when the number of positive samples (objects of interest) is much smaller than the number of negative samples (background). In such cases, the model tends to focus on the negative samples and neglect the positive samples, leading to poor performance. The focal loss addresses this issue by down-weighting the easy negative samples and up-weighting the hard positive samples.
The focal loss is defined as
$FL(p_{t})=-\alpha_{t}(1-p_{t})^{\gamma}log(p_{t}),$ (23)
where $p_{t}$ is the predicted probability for the true class, $\alpha_{t}$ is a weighting factor that controls the importance of each example, and $\gamma$ is a focusing parameter that controls the rate at which easy examples are down-weighted.
The weighting factor $\alpha_{t}$ is usually set to the inverse class frequency to balance the loss across all classes. The focusing parameter $\gamma$ is typically set to a value between 2 and 4 to give more weight to hard examples.
In the original paper, the authors used a sigmoid activation function for binary classification and the cross-entropy loss for multi-class classification. The focal loss is combined with these loss functions to improve the performance of object detection and semantic segmentation models.
In recent works, focal loss has been used in object detection, semantic [63], instance segmentation [64], and human pose estimation [65].
4.1.7 Hinge Loss
Hinge loss is a popular function used for maximum-margin classification, commonly used for support vector machines (SVMs) for example in one-vs-all classification where we classify an instance as belonging to one of many categories and situations where we want to provide a margin of error [66].
The hinge loss function for an individual instance can be represented as
@@ 4.1.7 Hinge Loss
where $y$ is the true label of the instance, which should be -1 or 1 in a binary classification problem. $f(x)$ is the predicted output for the instance $x$. The raw margin is $y\cdot f(x)$.
The hinge loss is 0 if the instance is on the correct side of the margin. The loss is proportional to the distance from the margin for data on the wrong side of the margin.
4.2 Classification Performance Metrics
Table 4 summarizes the common metrics used for classification. The following sections will delve into each of these metrics.
@@ 4.2.1 Confusion Matrix
For example, consider a binary classification problem where the algorithm tries to predict whether an image contains a cat. The confusion matrix for this problem would look like Table 6:
Table 6: Confusion Matrix
Predicted as Cat Predicted as Not Cat
Actual Cat TP FN
Actual Not Cat FP TN
Where:
  • TP: the number of images correctly classified as cats.
  • TN: the number of images correctly classified as not cats.
  • FP: the number of images incorrectly classified as cats.
  • FN: the number of images incorrectly classified as not cats.
@@
4.2.2 Accuracy
Accuracy is a commonly used metric for object classification. It is the ratio of correctly classified samples to the total number of samples [67]. Mathematically, it can be represented as
@@ 4.2.2 Accuracy
Accuracy can be expressed in terms of the confusion matrix values as
@@
It is a simple and intuitive metric, but it can be misleading when the class distribution is imbalanced, as it tends to favor the majority class. For example, let’s assume that we want to predict the presence of cancer in a cell. If for every 100 samples, only one contains cancer, a useless model that always predicts "No cancer" will have an accuracy of 99%. Other metrics, such as precision, recall, or F1-score, are more appropriate in these cases.
4.2.3 Precision
Precision measures the accuracy of positive predictions. It is defined as the number of true positive predictions divided by the number of true positive predictions plus the number of false positive predictions [68]. Mathematically, it can be represented as
$Precision=\frac{TP}{TP+FP},$ (27)
@@ 4.2.3 Precision
where $TP$ is the number of true positive predictions, and $FP$ is the number of false positive predictions.
Precision is useful when the cost of a false positive prediction is high, such as in medical diagnosis or fraud detection. A high precision means the model is not generating many false positives, so the predictions are reliable. However, it is important to note that precision is not the only metric to consider when evaluating a model’s performance, as high precision can also be achieved by a model that is not generating many positive predictions at all, which would result in a low recall.
4.2.4 Recall, Sensitivity, or True Positive Rate (TPR)
The recall metric, also known as sensitivity or True Positive Rate (TPR), measures the proportion of true positive instances (i.e., instances correctly classified as positive) out of the total number of positive instances [68]. Mathematically, it is represented as
$Recall=\frac{TruePositives}{TruePositives+FalseNegatives}$ (28)
It measures how well the model can identify all the positive instances in the dataset. A high recall value indicates the model has fewer false negatives, meaning it can correctly identify the most positive instances. However, a high recall value does not necessarily mean the model has a high precision, as the number of false positives can also influence it.
4.2.5 Precision-Recall Tradeoff
@@
The precision-recall tradeoff refers to the inverse relationship between precision and recall. As one metric increases, the other tends to decrease.
Imagine a machine learning model trying to predict whether an email is spam. If the model is tuned to be very conservative and only marks an email as spam when confident, it is likely to have high precision (i.e., if it marks an email as spam, it is very likely to be spam). However, this conservative approach means it will probably miss many spam emails it is unsure about, leading to a lower recall.
Conversely, if the model is tuned to be liberal and marks emails as spam more freely, it will probably identify most spam emails correctly, leading to a high recall. However, this approach will also incorrectly mark many non-spam emails as spam, leading to a lower precision.
@@
This tug-of-war between precision and recall is the crux of the tradeoff. An optimal balance between the two must be found depending on the use case. For instance, in a medical context, a high recall might be prioritized to ensure that all possible disease cases are identified, even at the expense of some false positives. On the other hand, a spam detection system might aim for high precision to avoid annoying users with wrongly classified emails, accepting that some spam messages might slip through.
The precision-recall tradeoff is a crucial consideration when tuning machine learning models. Maximizing both metrics is only sometimes possible; thus, a balance must be struck based on the requirements and constraints of the specific application.
4.2.6 F1-score
@@
The F1 score combines precision and recall to provide a single value representing a classification model’s overall performance [68]. It is defined as the harmonic mean of precision and recall computed as
$F1=2*\frac{precision\cdot recall}{precision+recall}$ (29)

### v3

@@ 3.2 Classification
Classification is a supervised machine learning task that trains a model to predict the category or class of a given input data point. The primary goal of classification is to help the model establish a mapping between input features and a specific class or category.
@@
There are different types of classification tasks, including binary classification, multi-class classification, and multi-label classification. Binary classification involves training the model to predict one of two classes, such as "spam" or "not spam," for an email. On the other hand, multi-class classification requires the model to predict one of several classes, such as "dog," "cat," or "bird," for an image. In multi-label classification, the model is trained to predict multiple labels for a single data point, such as "dog" and "outdoor," for an image of a dog in the park.
@@
Classification algorithms are based on various techniques such as decision trees, Naive Bayes, k-nearest neighbors, Support Vector Machines, Random Forest, Gradient Boosting, Neural Networks, and others. In the following subsections, we will delve deeper into the most common classification loss functions and performance metrics.
3.2.1 Classification Loss Functions
@@
Several loss functions can be used for classification tasks, depending on the specific problem and algorithm. Table 6 summarizes the most common loss functions used in classification. In the following sections, we describe each of these loss functions with more detail:
Table 6: Loss Functions and their applications in classification tasks.
Loss Function Applications
Binary Cross-Entropy (BCE) Binary classification, Logistic regression, Neural networks
Weighted Binary Cross-Entropy (WBCE) Imbalanced binary classification, Rare event prediction
Categorical Cross-Entropy (CCE) Multiclass classification, Deep learning models
Sparse Categorical Cross-Entropy Multiclass classification with large number of classes, Natural language processing
Cross-Entropy with Label Smoothing Multiclass classification with noisy labels, Neural networks to improve generalization
PolyLoss Imbalanced data, multiple tasks, 2D image classification, 2D and 3D object detection, instance segmentation.
Hinge Loss Support vector machines, Binary classification with margin maximization
Binary Cross-Entropy Loss (BCE)
@@
Binary Cross-Entropy Loss, also known as log loss, is a prevalent loss function for binary classification problems. It measures the dissimilarity between the predicted probability and the true class label [57]. Cross-entropy originates from information theory, where it measures the dissimilarity between two probability distributions.
In the context of binary classification, where the classes are labeled as 0 or 1, the BCE loss for an individual prediction is defined as:
@@
This formula can be intuitively understood by considering it in two parts:
@@
where $y$ is the true class label and $p$ is the predicted probability of the class being 1. The loss is minimized when the predicted probability $p$ matches the actual label $y$.
BCE is particularly useful because it is differentiable and provides a probabilistic interpretation of the model’s output, essential for gradient-based optimization algorithms. It also naturally penalizes incorrect classifications with high confidence, aligning well with the goals of logistic regression, which uses BCE as its loss function.
However, BCE can be sensitive to class imbalance, where the number of instances in one class significantly outnumbers the other. To address this, a variant called Weighted Binary Cross-Entropy can be used, which assigns a higher weight to the minority class, helping to balance the influence of each class on the training process.
$L_{weighted}(y,p)=-\left(w_{1}\cdot y\log(p)+w_{0}\cdot(1-y)\log(1-p)\right),$ (19)
where $w_{1}$ and $w_{0}$ are weights for the positive and negative classes, respectively. These weights are typically set inversely proportional to the class frequencies.
Categorical Cross-entropy Loss (CCE)
@@
Categorical Cross-Entropy Loss, also known as multi-class log loss, is used in multi-class classification tasks to measure the dissimilarity between the predicted probabilities and the actual distribution of the outcomes. It extends the concept of binary cross-entropy to multiple classes.
@@
The loss for categorical cross-entropy is defined as the average negative log-likelihood of the correct class:
$L=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{C}y_{i,j}\log(p_{i,j}),$ (20)
where $N$ is the number of samples, $C$ is the number of classes, $y_{i,j}$ is the true label in a one-hot encoded format, and $p_{i,j}$ is the predicted probability of class $j$ for sample $i$. This formulation means that the loss calculates the logarithm of the predicted probability for the actual class and penalizes the model based on the distance from the actual label.
One-hot encoding represents the true class of each sample with a vector of length $C$, where the index of the true class is marked with a 1, and all other entries are 0. This method ensures that the model’s prediction for the correct class is strongly reinforced if accurate and penalized if not.
CCE is critical for models where the accurate prediction of multiple classes is necessary. It is differentiable, making it suitable for optimization algorithms such as gradient descent used in training deep neural networks. Like binary cross-entropy, it severely penalizes incorrect confident predictions, which enhances the learning efficacy of classifiers.
Sparse Categorical Cross-entropy Loss
Sparse Categorical Cross-Entropy Loss is a variation of the standard categorical cross-entropy loss designed for multi-class classification tasks where the classes are encoded as integers rather than one-hot encoded vectors. This approach is particularly useful in scenarios with a large number of classes, where one-hot encoding would be memory-intensive and less efficient.
The loss for an individual prediction is calculated directly using the integer label as an index:
@@
The overall Sparse Categorical Cross-Entropy Loss for the dataset is the average of these individual losses:
@@
where $y_{i}$ is the integer label representing the true class of the $i$-th sample, and $\hat{y}_{i,y_{i}}$ is the predicted probability of that class. This formulation allows for a direct reference to the predicted probability corresponding to the actual class label, simplifying the computational process.
Sparse Categorical Cross-Entropy is differentiable, making it suitable for use with gradient descent and other optimization algorithms essential for training deep neural networks. Sparse CCE not only enhances computational efficiency but also aligns well with situations where the categorical labels are naturally ordinal or where the encoding of categories as integers simplifies data preprocessing and handling.
Cross-Entropy Loss with Label Smoothing
Cross-Entropy loss with label smoothing is an advanced regularization technique used to prevent models from becoming overly confident in their predictions by smoothing the labels. This technique modifies the standard Cross-Entropy loss by adjusting the target labels: a small constant $\epsilon$ is subtracted from the target class and evenly distributed among the other class labels. This approach is particularly effective in mitigating overconfidence, a common issue when training on large datasets, which can lead to degraded performance on unseen data [58, 59].
@@
The formula for Cross-Entropy loss with label smoothing is adjusted as follows:
$L(y,\hat{y})=-\sum_{c=1}^{C}\left[(1-\epsilon)y_{c}\log\hat{y}_{c}+\frac{\epsilon}{C-1}(1-y_{c})\log\hat{y}_{c}\right],$ (23)
where $y$ represents the true label vector, $\hat{y}$ the predicted probabilities, $C$ the number of classes, and $\epsilon$ a small constant (often set between 0.1 and 0.2). Here, each incorrect class receives a portion of $\epsilon$ divided by $C-1$, ensuring that the sum of probabilities remains normalized.
Label smoothing adjusts the confidence level of the predictions by penalizing extreme values in the log probabilities, thereby encouraging the model to allocate probability mass more evenly across different classes. This technique has been shown to improve the generalization of models, particularly in scenarios with many categories or when the dataset contains noisy labels. However, the optimal value of $\epsilon$ may vary depending on specific task requirements and characteristics of the data, often requiring empirical tuning to achieve the best results.
Negative Log-likelihood
In many machine learning applications, the true distribution $p$ is represented as a one-hot encoded vector, where only the true class has a probability of 1, and all others have 0. The predicted distribution $q$ is the output of the model. The cross-entropy loss for a single instance can be expressed as:
$H(p,q)=-\sum_{i}p_{i}\log(q_{i}),$ (24)
where $p$ is the true distribution and $q$ is the predicted distribution. Given that the true distribution $p$ is one-hot encoded, the summation simplifies to:
$H(p,q)=-\log(q_{y_{t}}),$ (25)
where $y_{t}$ denotes the true class. This expression is equivalent to the Negative Log Likelihood (NLL), which is defined as:
$\text{NLL}=-\log(p(y_{t}|x))$ (26)
Thus, in classification tasks with one-hot encoded labels, minimizing the cross-entropy loss is equivalent to minimizing the negative log-likelihood. Both metrics serve to encourage the model to assign a higher probability to the true class.
PolyLoss
PolyLoss [60] generalizes common loss functions like cross-entropy and focal loss. Represents loss functions as a linear combination of polynomial functions, offering a new perspective on how these functions can be adapted for specific tasks and datasets.
The PolyLoss framework is inspired by the Taylor series expansion, where the loss function $L$ is expressed as:
$L=\sum_{n=0}^{N}\alpha_{n}(1-p_{t})^{n},$ (27)
where $p_{t}$ is the predicted probability for the target class, and $\alpha_{n}$ are the coefficients for the $n$-th polynomial term. This framework provides a mechanism for adjusting the polynomial coefficients $\alpha_{n}$, which in turn modulates the sensitivity and emphasis of the loss function on various aspects of the predictions.
Poly-1 loss
A variant within this framework is the Poly-1 loss, which introduces a single additional hyperparameter $\epsilon$. The Poly-1 loss can be formulated as:
$L_{\text{Poly-1}}=\text{CE}+\epsilon(1-p_{t}),$ (28)
where CE is the standard cross-entropy loss, and $\epsilon$ is a hyperparameter that adjusts the influence of the additional term. By setting $\epsilon=0$, Poly-1 reduces to the standard cross-entropy loss. When $\epsilon>0$, the loss function becomes more sensitive to confident predictions, reducing overfitting in imbalanced datasets or tasks requiring higher precision.
PolyLoss is particularly beneficial in the following scenarios:
  1. 1.
    When dealing with imbalanced datasets, traditional loss functions like cross-entropy and focal loss may not perform optimally. PolyLoss can be tailored to adjust the importance of different classes through its polynomial coefficients, making it more effective for such cases.
  2. 2.
    When working on various tasks or datasets, PolyLoss allows to modify polynomial coefficients for a customized approach, which can lead to better performance across different applications.
  3. 3.
    To simplify the hyperparameter optimization process. The Poly-1 formulation, for example, introduces only one additional hyperparameter, making it easier to tune compared to more complex loss functions.
Hinge Loss
Hinge Loss is a widely used loss function in maximum-margin classification tasks, particularly with Support Vector Machines (SVMs). It’s especially prevalent in scenarios requiring a robust classification margin, such as in one-vs-all classification tasks where an instance is classified into one of several categories [61].
The Hinge Loss for a given instance is formulated as:
@@
where $y$ represents the true label of the instance, taking values -1 or 1 in binary classification contexts, and $f(x)$ is the predicted output from the model for the input $x$. The term $y\cdot f(x)$ reflects the raw margin, which measures how far the predicted value is from the decision boundary in terms of alignment and distance.
The essence of Hinge Loss is that it is zero if the predicted classification is correct and the instance is beyond the margin (the correct side). If the classification is wrong, or if the classification is correct but too close to the decision boundary (the wrong side of the margin), the loss increases linearly with the distance from the margin.
Hinge Loss not only penalizes classification errors but also correct classifications that are not confidently made (i.e., those close to the decision boundary). This characteristic helps in forming a decision boundary that is not only accurate in separating classes but also maximizes the distance (margin) between the nearest points of the classes and the boundary itself, enhancing model generalization.
In addition to its basic form, variants such as the squared hinge loss ($\max(0,1-y\cdot f(x))^{2}$) are used to penalize incorrect classifications more strongly, promoting an even more robust classification boundary. Although primarily associated with binary classification, Hinge Loss can be adapted for multi-class problems through approaches like structured SVMs [62], which handle complex inter-class relationships.
3.2.2 Classification Performance Metrics
In this section, we will discuss the various metrics that are commonly used to evaluate the performance of classification models. Table 7 provides a summary of these metrics, and the following subsections describe each one in more detail.
@@ 3.2.2 Classification Performance Metrics
  1. 1.
    True Positives (TP): These are the cases where the model correctly predicts the positive class. In other words, the instances that are actually positive and are correctly identified as positive by the model.
  2. 2.
    True Negatives (TN): These are the cases where the model correctly predicts the negative class. This means the instances are actually negative and are identified as negative by the model.
  3. 3.
    False Positives (FP): Also known as Type I errors, these occur when the model incorrectly predicts the positive class. That is, the instances are actually negative but are wrongly labeled as positive. This is sometimes called "false alarm".
  4. 4.
    False Negatives (FN): Also known as Type II errors, these occur when the model incorrectly predicts the negative class. These instances are actually positive but are mistakenly labeled as negative. This is sometimes referred to as a "miss".
These metrics provide insights into how well the model performs and where it might be making the most mistakes, enabling targeted improvements.
@@
Confusion Matrix in Multi-class Classification
In multiclass classification, the confusion matrix is an extension of the binary classification confusion matrix, where there are more than two classes to predict. The matrix helps in visualizing the performance of the classification algorithm across all classes.
For a classification task with $N$ classes, the confusion matrix will be an $N\times N$ matrix. Each row of the matrix corresponds to the true class, while each column corresponds to the predicted class. The element at the $i$-th row and $j$-th column of the matrix, denoted as $M[i,j]$, represents the number of instances where the true class is $i$ but the predicted class is $j$.
The diagonal elements ($M[i,i]$) represent the number of points that were correctly classified as class $i$. Therefore, for all $i$ where $1\leq i\leq N$, $M[i,i]$ are the true positives for class $i$.
The off-diagonal elements ($M[i,j]$) where $i\neq j$) indicate misclassifications. Specifically, $M[i,j]$ is the number of observations that belong to class $i$ (true class) but were predicted as class $j$ (incorrect class).
For example, if we have three classes, A, B, and C, the confusion matrix might look something like Table 9.
Table 9: Confusion Matrix for Multiclass Classification. The diagonal elements $M[i,i]$ represent the number of points that were correctly classified as class $i$. The off-diagonal elements indicate misclassifications. Specifically, $M[i,j]$ is the number of observations that belong to class $i$ (true class) but were predicted as class $j$
Predicted: A Predicted: B Predicted: C
Actual: A $M[1,1]$ $M[1,2]$ $M[1,3]$
Actual: B $M[2,1]$ $M[2,2]$ $M[2,3]$
Actual: C $M[3,1]$ $M[3,2]$ $M[3,3]$
Multi-class Confusion Matrix Visualization
Multi-class confusion matrices are often presented as heatmaps as shown in Figure 1. This visualization represents the confusion matrix of a multi-class classification model with three classes: Class A, Class B, and Class C. Each cell in the matrix shows the number of samples from the true class (rows) that were predicted as each class (columns). The diagonal cells (from the top left to the bottom right) indicate the number of correct predictions for each class, highlighting the model’s accuracy for those specific categories.
Class A: Out of the total samples belonging to Class A, 95 were correctly predicted as Class A, while 2 were misclassified as Class B, and 3 as Class C.
Class B: For Class B, 90 samples were correctly classified, 4 were incorrectly predicted as Class A, and 6 as Class C.
Class C: In the case of Class C, 85 samples were correctly identified, with misclassifications including 7 as Class A and 8 as Class B. The non-diagonal cells reveal the instances of misclassification. For example, more samples of Class C were misclassified as Class B than as Class A, indicating specific areas where the model’s classification performance could be improved.
Figure 1: Confusion Matrix Heatmap for Multi-Class Classification. The matrix visualizes the performance of a classification model across three classes: Class A, Class B, and Class C. The values along the diagonal represent the number of correct predictions for each class, whereas off-diagonal values indicate misclassifications. The color intensity in each cell correlates with the number of samples, providing a clear visual distinction between higher and lower frequencies of predictions. This figure aids in identifying the strengths and weaknesses of the model, particularly in distinguishing between the classes..
Accuracy
Accuracy is a widely used performance metric in classification tasks, defined as the ratio of correctly classified samples to the total number of samples [63]. For a simple formula representation:
@@
In terms of a confusion matrix, which is a table used to describe the performance of a classification model, accuracy is calculated as:
@@
where TP, TN, FP, and FN represent the true positives, true negatives, false positives, and false negatives, respectively.
While straightforward and intuitive, accuracy can be misleading in datasets with imbalanced class distributions. For instance, in a dataset where 99 out of 100 samples are negative for a condition (e.g., no cancer), a model predicting ’No cancer’ for all instances would achieve 99% accuracy, despite not being useful in practical terms.
Accuracy in Multi-Class Classification
In multi-class scenarios, accuracy is extended by summing the correctly predicted instances across all classes and dividing by the total number of samples:
$Accuracy=\frac{\sum_{i=1}^{C}Correct_{i}}{Total\ Samples}$ (32)
where $C$ is the number of classes, and $Correct_{i}$ is the number of correctly predicted instances for each class. This measure treats all classes equally, but it does not account for the imbalance among different classes.
Balanced Accuracy
To address class imbalance, particularly in multi-class settings, balanced accuracy can be considered. This metric calculates the average of recall obtained on each class:
$Balanced\ Accuracy=\frac{1}{C}\sum_{i=1}^{C}\frac{TP_{i}}{TP_{i}+FN_{i}}$ (33)
where $TP_{i}$ and $FN_{i}$ are the true positives and false negatives for each class $i$. Balanced accuracy thus provides a more equitable measure of model performance across all classes.
Precision, PPV, or TDR
Precision, also known as positive predictive value (PPV) or true discovery rate (TDR), measures the accuracy of positive predictions made by a model. It is defined as the number of true positive predictions divided by the total number of predicted positives, which includes both true positives and false positives [64]. The formula for precision can be expressed as:
$\text{Precision}=\frac{TP}{TP+FP},$ (34)
@@
where $TP$ represents the number of true positive predictions, and $FP$ stands for the number of false positive predictions.
Precision is particularly critical in domains where the consequences of false positives are significant, such as in medical diagnostics or fraud detection. High precision indicates that the model produces few false positives, leading to more reliable predictions. However, it’s crucial to consider precision in conjunction with other performance metrics like recall, as a model could achieve high precision at the cost of recall, thereby missing a substantial number of actual positives.
Multi-Class Precision
In multi-class classification scenarios, precision is calculated for each class individually and then averaged, which can be done in several ways depending on the context and specific requirements of the evaluation. The two common methods are:
  • Macro-average Precision: This method calculates precision independently for each class and then takes the average. This approach treats all classes equally, regardless of their support (frequency). Mathematically, it is defined as:
[…]
    where $N$ is the number of classes, and $TP_{i}$ and $FP_{i}$ are the true positives and false positives for the $i$-th class, respectively.
  • Micro-average Precision: This method aggregates the contributions of all classes to compute the average precision. It calculates the total true positives and false positives across all classes and then applies the precision formula. This approach is influenced more by the class with a larger number of instances. It can be represented as:
[…]
    where $N$ is the number of classes.
Both averaging methods are useful but serve different purposes depending on the objective of the analysis. Macro-average is beneficial when you want to treat all classes equally, while micro-average is preferable when you expect a balance between classes to be reflected in the precision metric.
Recall, Sensitivity, or True Positive Rate (TPR)
Recall, also known as sensitivity or true positive rate (TPR), is a measure of a model’s ability to correctly identify positive instances among all actual positives in the dataset. It is calculated as the ratio of true positives to the sum of true positives and false negatives, where a true positive is an instance correctly identified as positive, and a false negative is a positive instance that was incorrectly identified as negative [64]. Mathematically, recall can be expressed as:
$\text{Recall}=\frac{TP}{TP+FN},$ (37)
where $TP$ is the number of true positives, and $FN$ is the number of false negatives.
A high recall score indicates that the model is successful in capturing a large proportion of positive cases, which is especially important in applications where failing to detect positives (such as diseases or fraudulent transactions) could have severe consequences. However, optimizing for recall may increase the number of false positives, thereby reducing precision. It is common practice to balance recall with precision and other metrics based on the specific needs of the application.
Multi-Class Recall
In multi-class classification problems, recall must be assessed for each class and then averaged, depending on the specific needs and the nature of the dataset. There are generally two methods to average recall in multi-class settings:
  • Macro-average Recall: This method calculates recall for each class independently and then takes the average of these measures. This treats all classes equally, regardless of their prevalence. The formula for macro-average recall is:
[…]
    where $N$ is the number of classes, $TP_{i}$ is the true positives for the $i$-th class, and $FN_{i}$ is the false negatives for the $i$-th class.
  • Micro-average Recall: This method aggregates the contributions of all classes to compute the overall recall. It is particularly useful when class imbalance is significant. The total number of true positives and false negatives across all classes are summed up, and the recall formula is applied to these totals:
[…]
    where $N$ is the number of classes.
Both averaging techniques are valuable and can be selected based on whether class equality or dataset representation is more important for the analytical goal.
Precision-Recall Tradeoff
@@
The precision-recall tradeoff describes the inverse relationship between precision and recall. As one metric increases, the other tends to decrease. This inverse relationship is crucial in settings where decisions need to balance the risks of false positives and false negatives.
Consider a machine learning model used for spam detection. If the model is set to be very conservative (high threshold), it is likely to mark only the most obvious spam emails as spam. This approach tends to result in high precision because most emails classified as spam are truly spam, but it may suffer from low recall as some spam emails are missed. Conversely, if the model is set to be more liberal (low threshold), it marks more emails as spam. This method increases recall by identifying most spam emails but at the expense of precision, as more legitimate emails are incorrectly flagged as spam.
@@
This tradeoff between precision and recall is fundamental in many applications. For instance, in medical diagnostics, a high recall might be prioritized to ensure that all potential disease cases are identified, even at the cost of some false positives. On the other hand, in spam detection systems, high precision might be more desirable to prevent legitimate emails from being incorrectly marked as spam, even if some spam emails pass through the filters.
The balance between precision and recall can be illustrated through a Precision-Recall curve (refer to section 3.2.2). This curve shows how these two metrics change based on different thresholds. The F1 score is a commonly used metric for situations where a single metric is required to balance precision and recall without prioritizing either one. We will discuss this metric in more detail in the following section.
F1-score
@@
The F1 score is a statistical measure that combines precision and recall to provide a single value representing a classification model’s overall performance [64]. It is defined as the harmonic mean of precision and recall, computed as:
$F1=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (40)